# Bosenovae with Quadratically-Coupled Scalars in Quantum Sensing Experiments

### v1

@@ III Boson Stars and Bosenovae
assuming attractive self-interactions $\lambda>0$. Subsequently, the boson star collapses. In this work, we are interested in transient signals associated with such events.
@@
Key properties of the bosenova emission from a boson star have been obtained with numerical simulations Levkov:2016rkk . Of interest for our work, the leading peak of the emission spectrum occurs around particle momentum of $\bar{k}\simeq 2.4m_{\phi}$ with a spread of $\delta k\simeq m_{\phi}$ (see also Ref. Eby:2015hyx ; Eby:2017azn ). The total energy emitted in this range can be estimated as Eby:2021ece ; Arakawa:2023gyq
@@
After emission, the relativistic ULDM wave will spread in-flight along the radial direction from the source. We briefly outline the consequences of these effects below, with a comprehensive discussion given in Ref. Eby:2021ece ; Arakawa:2023gyq . For emission sources located at astrophysical distances $r$, the wave spreading spreading in-flight rapidly comes to dominate the intrinsic burst duration $\delta t_{0}$. This can result in a persistent signal of a significant duration at detector site
@@
The expected rate of bosenovae in the Galaxy is a complex topic worthy of a dedicated study beyond the scope of the present work. The predictions depend on variety of considerations including cosmological evolution Gorghetto:2018myk ; Vaquero:2018tib ; Buschmann:2019icd ; Gorghetto:2020qws ; Buschmann:2021sdq ; Saikawa:2024bta and structure formation Schive:2014hza ; Schive:2014dra ; Mocz:2017wlg ; Schwabe:2020eac ; Nori:2020jzx for ULDM, as well as details of boson star formation Levkov:2018kau ; Chen:2020cef ; Kirkpatrick:2020fwd ; Chen:2021oot ; Kirkpatrick:2021wwz , tidal stripping Dokuchaev:2017psd ; Kavanagh:2020gcy ; Edwards:2020afl , accretion Eggemeier:2019jsu ; Chan:2022bkz ; Dmitriev:2023ipv and mergers Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh , among others. See also Refs. Eby:2021ece ; Arakawa:2023gyq for further discussion.
IV Detection
IV.1 Detection methods
A variety of distinct experiments can sensitively probe variations in fundamental constants that can result from ULDM interactions with SM, including the fine-structure constant $\alpha$. We refer to Ref. Antypas:2022asj for detailed discussion of experimental probes, and below summarize the searches which are relevant to our results discussed in Sec. V.
@@ IV.1 Detection methods
    Limits to $d_{m_{e}}$ include ${}^{133}$s/${}^{87}$b atomic fountain clock frequency ratio Hees:2016gop , frequency comparison between ${}^{171}$b optical lattice clock and ${}^{133}$s fountain microwave clock 2022PhRvL.129x1301K , and a comparison of Rb hyperfine transition with quartz oscillator Zhang:2022ewz .
    The limits on $d_{e}$ come from frequency ratio measurements of ${}^{27}$${}^{+}$, ${}^{171}$b and ${}^{87}$r optical clocks (BACON) Beloy:2020tgz , frequency ratios of the E3 transition in a single-ion ${}^{171}$${}^{+}$ clock with E2 transition in the same ion and with and ${}^{171}$${}^{+}$ ${}^{87}$r optical lattice clock Filzinger:2023zrs , and a dynamical decoupling demonstration in ${}^{+}$ trapped ion clock transition Aharony:2019iad .
    The limits on $d_{g}$ come from the ${}^{133}$s/${}^{87}$b atomic fountain clock frequency ratio Hees:2016gop , the limits from oscillating nuclear radius effects in ${}^{171}$${}^{+}$ ion clock Banerjee:2023bjc , and ${}^{171}$b optical lattice clock and ${}^{133}$s fountain microwave clock comparison 2022PhRvL.129x1301K . A significant improvement in $d_{m_{e}}$ limits is expected with the SrOH molecular experiment SrOH2021 . Very high sensitivity to variation of fundamental constants is expected of nuclear clock 2021Peik , which rely on a nuclear transition in ${}^{229}$h, leading to potential orders of magnitude improvement in $d_{e}$ and $d_{g}$ Antypas:2022asj .
@@
  • Spectroscopy: Variations in fundamental constants due to ULDM-SM couplings that lead to observable signatures in atomic and molecular spectra are the basis of clock searches. However, other spectroscopy experiments can also be used for this purpose. These include molecular spectroscopy of iodine molecule Oswald:2021vtc , frequency comparisons of ${}^{164}$y with quartz oscillator Zhang:2022ewz , and between two isotopes of dysprosium (Dy/Dy) VanTilburg:2015oza , which are all sensitive to $d_{e}$.
  • Mechanical resonators: Both torsion balances and optical-cavity-based UDM searches involve the measurement of the deformation of an elastic body (the mechanical response) produced by a weak force Antypas:2022asj . The deformation can be amplified by a factor as large as the resonance Q factor if the body has internal resonances at the ULDM Compton frequency, yielding massive sensitivity enhancement over a narrow bandwidth.
@@
We refer to the searches above as direct DM (DDM) searches due to the nature of their coupling directly to the DM background (see Sec. IV.2 for details). In addition to these direct searches, there are searches for ULDM emission or absorption in physical systems, which are outlined below. As we will explain in Sec. IV.2, the sensitivity of such probes will scale differently with the local DM density compared to direct searches.
@@
  • Fifth-force tests: Scalar coupling to SM fields can lead to Yukawa-type interaction between matter, modifying Newton’s law of universal gravitation. Tests of such “fifth forces” can constrain linear ULDM-SM couplings Fischbach:1996eq ; Murata:2014nra .
  • Astrophysical bounds: Massive self-interacting ultralight fields contribute to the evolution of the matter power spectrum (see our previous work for more details Arakawa:2023gyq ). We have outlined several astrophysical constraints on the ULDM parameter space in Sec. II.3. Constraints on photon emissivity in supernova cores from astrophysical observations can also be used to constrain ULDM-photon coupling Olive:2007aj .
Figure 4: The sensitivity ratio of a given DM-SM coupling for a burst search $d_{i,*}$ compared to a DM search $d_{i}$, using Eq. (26), is depicted in the allowed parameter space $m_{\phi}$ vs $\lambda$. The white dotted line represents equal sensitivity to bosenova and DM. Two choices for the distance to the bosenova are illustrated: $r=\rm{pc}$ (left) and $r=\rm{kpc}$ (right). See Sec. II.3 and Fig. 3 for a discussion of constraints.
IV.2 Differences with linear couplings
@@
Finally, the quadratic coupling case can induce screening of the field (as discussed in Sec. II), which is not the case for linear couplings. The terrestrial experiments are therefore at a disadvantage because they cannot probe arbitrarily large dilatonic couplings, and must reach sensitivities beyond the critical dilatonic coupling in Eq. (6). All direct-detection and EP experiments that are based on the Earth are subject to this screening, which provides important motivation for future space-based searches.
V Results
@@
We find that bosenovae can lead to large enhancements of the ULDM density at experimental site in a wide range of the parameter space. Our results complement those of Ref. Arakawa:2023gyq that analyzed bosenovae signatures for linear ULDM couplings to SM constituents. However, in our studied case of quadratic couplings, the signal enhancement is greater than for linear couplings. This is because the sensitivity reach of quadratic couplings follows a scaling relation of $d^{(2)}_{i,*}/d^{(2)}_{i,{\rm DM}}\sim\rho_{\rm DM}/\rho_{*}$, as opposed to $d^{(1)}_{i,*}/d^{(1)}_{i,{\rm DM}}\sim\sqrt{\rho_{\rm DM}/\rho_{*}}$ for linear couplings Arakawa:2023gyq . We demonstrate the sensitivity enhancement in the case of a bosenova search using Eq. (26) for the parameter space of $m_{\phi}-\lambda$ plane in Fig. 4. The region below the dotted white line indicates $d_{i,*}^{(2)}/d_{i,{\rm DM}}^{(2)}<1$, implying improved sensitivity to a bosenova search provided that such a transient event occurs within distance $r$ from detector. We find that the reach for ULDM dilatonic couplings could be over 10 orders of magnitude better than that of cold DM in case of a bosenova occurring below $\lesssim$ kpc of experimental site.
@@
In Figs. 5, 6, and 7, we display the electron ($d^{(2)}_{m_{e}}$), photon ($d_{e}^{(2)}$), and gluon ($d_{g}^{(2)}$) ULDM coupling reach associated with bosenova events as a function of $E_{\phi}=2.6\,m_{\phi}$, where the energy of the scalar field is identified with the expected from simulations leading peak in the emission spectrum from a boson star explosion Levkov:2016rkk . We showcase results for two benchmark distances between the Earth and the bosenova of $r=1~{}{\rm pc}$ (dashed black) and $r=1~{}{\rm kpc}$ (dotted black), considering reference coupling $\lambda$ from Eq. (12). Each figure displays four panels, separated by the current (left column) and the projected (right column) experiments, and terrestrial (top row) as well as space-based (bottom row) experiments. The terrestrial experiments are subject to screening due to the Earth, demonstrated by the light blue hashed critical screening region, while the space-based experiments are not. Since the screening by Earth creates an upper bound on dilatonic couplings able to be probed by terrestrial experiments, we cut off the black lines if they continue above the screening region. We also represent the region where BHSR bounds would affect the results, without taking them into account for our benchmark (see Sec.  II.3). We discuss the results with the BHSR constraints included in App. A.
@@
We note that the astrophysical constraints coming from Lyman-$\alpha$ and ultra-faint dwarf galaxies (UFDs) described in Sec. II.3 rely on the assumption that $\phi$ comprises the entirety of the DM abundance. Hence, they become weaker or even disappear if $\phi$ represents a sub-leading contribution to the total DM density. The laboratory bounds therefore give complementary constraints on such mass regions.
@@
Figure 7: Current and projected bounds on $d^{(2)}_{g}$ from terrestrial and space-based experiments as a function of $E_{\phi}$. The best bounds for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. We choose $\lambda=\lambda_{b}$ in Eq. (12), as explained in the text. The terrestrial experiments contributing to current bounds on $d_{g}^{(2)}$ include Clocks: ${}^{171}$b lattice clock - ${}^{133}$s microwave clock comparison (Yb/Cs) 2022PhRvL.129x1301K , frequency comparison in dual rubidium-caesium cold atom clock (Rb/Cs) Hees:2016gop and ${}^{+}$ ion clock Banerjee:2023bjc ; and Optical cavities: H-maser comparison with a Si cavity (H/Si) Kennedy:2020bac ; while space-based current bound comes from Equivalence-principle tests: MICROSCOPE Berge:2017ovy . Future experiment projections are for the ${}^{229}$h nNuclear cClock Antypas:2022asj and ${}^{171}$${}^{+}$ ion clock Banerjee:2023bjc . Note: The Nuclear Clock line in previous estimates is cut off at $10^{-21}~{}\rm eV$, due to the Lyman-$\alpha$ constraint. However, there is no reason that the experiment cannot be run to search for lower mass range, as in $\rm Yb^{+}$ experiment. Therefore, we extrapolate linearly to lower masses. We are not aware of any space-based planned experiments that will constrain $d^{(2)}_{g}$. In cyan, we show the region of the parameter space critically screened by the Earth, as discussed in Sec. II.2. Lyman-$\alpha$ Irsic:2017yje ; Rogers:2020ltq and ultra-faint dwarf galaxies (UFDs) Dalal:2022rmp also constrain ULDM. Finally, the regions of the parameter space corresponding to the black hole superradiance (BHSR) bounds Baryakhtar:2020gao ; Unal:2020jiy are displayed at the bottom of the plots by the arrows. In the terrestrial plots, we also show the region of parameter space critically screened by the Earth in cyan. See Sec. II.2 for a detailed discussion of screening of quadratically-coupled scalars.
VI Conclusions

### v2

@@ 3 Boson Stars and Bosenovae
assuming attractive self-interactions $\lambda>0$. Subsequently, the boson star collapses. Due to the minus sign on the one-loop electron correction (see Fig. 2 and Eq. 13), one might consider whether or not the electron loops could contribute more during the final moments of the boson star collapse, and cause the $\lambda$ to flip sign, resulting in a respulsive self-interaction. However, since the energies associated with the bosenova are $\sim$ a few×$\,\times\,m_{\phi}$, the shift in energy of the process is only a factor of a few, which is still much lower than the electron mass. We therefore do not expect this to be an important effect. In this work, we are interested in transient signals associated with such events.
@@
Key properties of the bosenova emission from a boson star have been obtained with numerical simulations Levkov:2016rkk . Of interest for our work, the leading peak of the emission spectrum occurs around particle momentum of $\bar{k}\simeq 2.4m_{\phi}$ with a spread of $\delta k\simeq m_{\phi}$ (see also Ref. Eby:2015hyx ; Eby:2017azn ). This first peak can be understood by energy conservation, as it is expected to be generated by 3-to-1 ($\phi\phi\phi\rightarrow\phi$) annihilation, where the initial state consists of bound state $\phi$ with energies comparable to their masses. Energy conservation then implies an energy for the outgoing $\phi$ of order $\sim 3m_{\phi}$. The total energy emitted in this range can be estimated as Eby:2021ece ; Arakawa:2023gyq
@@
After emission, the relativistic ULDM wave will spread in-flight along the radial direction from the source. We briefly outline the consequences of these effects below, with a comprehensive discussion given in Ref. Eby:2021ece ; Arakawa:2023gyq . For emission sources located at astrophysical distances $r$, the wave spreading spreading in-flight rapidly comes to dominate the intrinsic burst duration $\delta t_{0}\simeq 400/m_{\phi}$ Levkov:2016rkk , which is intuitively proportional to the characteristic oscillation timescale of the bosons, $1/m_{\phi}$. However, we find that given the relatively low energies of the emitted bosons, $E_{\phi}\simeq{\rm few}\times m_{\phi}$, the effect of wave-spreading dominates the signal duration, and the internal burst time does not significantly modify the result. Due to this, we only need to consider the signal duration due to wavespreading. This can result in a persistent signal of a significant duration at detector site
@@
The expected rate of bosenovae in the Galaxy is a complex topic worthy of a dedicated study beyond the scope of the present work. The predictions depend on variety of considerations including cosmological evolution Gorghetto:2018myk ; Vaquero:2018tib ; Buschmann:2019icd ; Gorghetto:2020qws ; Buschmann:2021sdq ; Saikawa:2024bta and structure formation Schive:2014hza ; Schive:2014dra ; Mocz:2017wlg ; Schwabe:2020eac ; Nori:2020jzx for ULDM, as well as details of boson star formation Levkov:2018kau ; Chen:2020cef ; Kirkpatrick:2020fwd ; Chen:2021oot ; Kirkpatrick:2021wwz , tidal stripping Dokuchaev:2017psd ; Kavanagh:2020gcy ; Edwards:2020afl , accretion Eggemeier:2019jsu ; Chan:2022bkz ; Dmitriev:2023ipv and mergers Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh , among others. See also Refs. Eby:2021ece ; Arakawa:2023gyq for further discussion. Recent studies, under various simplifying assumptions, have estimated the DM fraction expected to consist of boson stars finding fractions of few to tens of percent Gorghetto:2024vnp ; Chang:2024fol . Additionally, Ref. Maseizik:2024qly found that the rate of bosenovae in a Milky Way-like galaxy can be as large as a few per day, depending on the model. Given the uncertainties and complexity of population statistics and event rate estimates, in this study, instead, we focus on characterizing signals and features of a singular bosenova event.
We emphasize that although we focus on bosenovae events here, our analysis methodology is applicable to other potential sources of transient signals, such as instabilities of superradiant clouds around black holes or gravitational atoms around stars. Recently, it has been shown that intriguing signatures can result from accumulation of such emissions over time to comprise, in case of axions, diffuse axion background Eby:2024mhd .
4 Detection
4.1 Detection methods
A variety of distinct experiments can sensitively probe variations in fundamental constants that can result from ULDM interactions with SM, including the fine-structure constant $\alpha$. We refer to Ref. Antypas:2022asj for detailed discussion of experimental probes, and below summarize the searches which are relevant to our results discussed in Sec. 5.
@@ 4.1 Detection methods
    Limits to $d_{m_{e}}$ include 133Cs/87Rb atomic fountain clock frequency ratio Hees:2016gop , frequency comparison between 171Yb optical lattice clock and 133Cs fountain microwave clock 2022PhRvL.129x1301K , and a comparison of Rb hyperfine transition with quartz oscillator Zhang:2022ewz .
    The limits on $d_{e}$ come from frequency ratio measurements of 27Al+, 171Yb and 87Sr optical clocks (BACON) Beloy:2020tgz , frequency ratios of the E3 transition in a single-ion 171Yb+ clock with E2 transition in the same ion and with and 171Yb+ 87Sr optical lattice clock Filzinger:2023zrs , and a dynamical decoupling demonstration in Sr+ trapped ion clock transition Aharony:2019iad .
    The limits on $d_{g}$ come from the 133Cs/87Rb atomic fountain clock frequency ratio Hees:2016gop , the limits from oscillating nuclear radius effects in 171Yb+ ion clock Banerjee:2023bjc , and 171Yb optical lattice clock and 133Cs fountain microwave clock comparison 2022PhRvL.129x1301K . A significant improvement in $d_{m_{e}}$ limits is expected with the SrOH molecular experiment SrOH2021 . Very high sensitivity to variation of fundamental constants is expected of the nuclear clock 2021Peik , which rely on a nuclear transition in 229Th, leading to potential orders of magnitude improvement in $d_{e}$ and $d_{g}$ Antypas:2022asj .
@@
  • Spectroscopy: Variations in fundamental constants due to ULDM-SM couplings that lead to observable signatures in atomic and molecular spectra are the basis of clock searches. However, other spectroscopy experiments can also be used for this purpose. These include molecular spectroscopy of iodine molecules Oswald:2021vtc , frequency comparisons of 164Dy with quartz oscillator Zhang:2022ewz , and between two isotopes of dysprosium (Dy/Dy) VanTilburg:2015oza , which are all sensitive to $d_{e}$.
  • Mechanical resonators: Both torsion balances and optical-cavity-based ULDM searches involve the measurement of the deformation of an elastic body (the mechanical response) produced by a weak force Antypas:2022asj . The deformation can be amplified by a factor as large as the resonance Q factor if the body has internal resonances at the ULDM Compton frequency, yielding massive sensitivity enhancement over a narrow bandwidth.
@@
We refer to the searches above as direct DM (DDM) searches due to the nature of their coupling directly to the DM background (see Sec. 4.2 for details). In addition to these direct searches, there are searches for ULDM emission or absorption in physical systems, which are outlined below. As we will explain in Sec. 4.2, the sensitivity of such probes will scale differently with the local DM density compared to direct searches.
@@
  • Fifth-force tests: Scalar coupling to SM fields can lead to Yukawa-type interactions between matter, modifying Newton’s law of universal gravitation. Tests of such “fifth forces” can constrain linear ULDM-SM couplings Fischbach:1996eq ; Murata:2014nra .
  • Astrophysical bounds: Massive self-interacting ultralight fields contribute to the evolution of the matter power spectrum (see our previous work for more details Arakawa:2023gyq ). We have outlined several astrophysical constraints on the ULDM parameter space in Sec. 2.3. Constraints on photon emissivity in supernova cores from astrophysical observations can also be used to constrain ULDM-photon coupling Olive:2007aj .
Figure 4: The sensitivity ratio of a given DM-SM coupling for a burst search $d_{i,*}$ compared to a DM search $d_{i}$, using Eq. (26), is depicted in the allowed parameter space $m_{\phi}$ vs $\lambda$. The white dotted line represents equal sensitivity to bosenova and DM. Two choices for the distance to the bosenova are illustrated: $r=\rm{pc}$ (left) and $r=\rm{kpc}$ (right). See Sec. 2.3 and Fig. 3 for a discussion of constraints.
4.2 Differences with linear couplings
@@
Finally, the quadratic coupling case can induce screening of the field (as discussed in Sec. 2), which is not the case for linear couplings. The terrestrial experiments are therefore at a disadvantage because they cannot probe arbitrarily large dilatonic couplings, and must reach sensitivities beyond the critical dilatonic coupling in Eq. (6). All direct-detection and EP experiments that are based on the Earth are subject to this screening, which provides important motivation for future space-based searches.
5 Results
@@
We find that bosenovae can lead to large enhancements of the ULDM density at experimental sites in a wide range of the parameter space. Our results complement those of Ref. Arakawa:2023gyq that analyzed bosenovae signatures for linear ULDM couplings to SM constituents. However, in our studied case of quadratic couplings, the signal enhancement is greater than for linear couplings. This is because the sensitivity reach of quadratic couplings follows a scaling relation of $d^{(2)}_{i,*}/d^{(2)}_{i,{\rm DM}}\sim\rho_{\rm DM}/\rho_{*}$, as opposed to $d^{(1)}_{i,*}/d^{(1)}_{i,{\rm DM}}\sim\sqrt{\rho_{\rm DM}/\rho_{*}}$ for linear couplings Arakawa:2023gyq . We demonstrate the sensitivity enhancement in the case of a bosenova search using Eq. (26) for the parameter space of $m_{\phi}-\lambda$ plane in Fig. 4. The region below the dotted white line indicates $d_{i,*}^{(2)}/d_{i,{\rm DM}}^{(2)}<1$, implying improved sensitivity to a bosenova search provided that such a transient event occurs within distance $r$ from detector. We find that the reach for ULDM dilatonic couplings could be over 10 orders of magnitude better than that of cold DM in case of a bosenova occurring below $\lesssim$ kpc of experimental site.
@@
In Figs. 5, 6, and 7, we display the electron ($d^{(2)}_{m_{e}}$), photon ($d_{e}^{(2)}$), and gluon ($d_{g}^{(2)}$) ULDM coupling reach associated with bosenova events as a function of $E_{\phi}=2.6\,m_{\phi}$, where the energy of the scalar field is identified with the expected from simulations leading peak in the emission spectrum from a boson star explosion Levkov:2016rkk . We showcase results for two benchmark distances between the Earth and the bosenova of $r=1~{}{\rm pc}$ (dashed black) and $r=1~{}{\rm kpc}$ (dotted black), considering reference coupling $\lambda$ from Eq. (12). Each figure displays four panels, separated by the current (left column) and the projected (right column) experiments, and terrestrial (top row) as well as space-based (bottom row) experiments. The terrestrial experiments are subject to screening due to the Earth, demonstrated by the light blue hashed critical screening region, while the space-based experiments are not. Since the screening by Earth creates an upper bound on dilatonic couplings able to be probed by terrestrial experiments, we cut off the black lines if they continue above the screening region. We also represent the region where BHSR bounds would affect the results, without taking them into account for our benchmark (see Sec.  2.3). We discuss the results with the BHSR constraints included in App. A.
@@
We note that the astrophysical constraints coming from Lyman-$\alpha$ and ultra-faint dwarf galaxies (UFDs) described in Sec. 2.3 rely on the assumption that $\phi$ comprises the entirety of the DM abundance. Hence, they become weaker or even disappear if $\phi$ represents a sub-leading contribution to the total DM density. The laboratory bounds therefore give complementary constraints on such mass regions.
@@
Figure 7: Current and projected bounds on $d^{(2)}_{g}$ from terrestrial and space-based experiments as a function of $E_{\phi}$. The best bounds for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. We choose $\lambda=\lambda_{b}$ in Eq. (12), as explained in the text. The terrestrial experiments contributing to current bounds on $d_{g}^{(2)}$ include Clocks: 171Yb lattice clock - 133Cs microwave clock comparison (Yb/Cs) 2022PhRvL.129x1301K , frequency comparison in dual rubidium-caesium cold atom clock (Rb/Cs) Hees:2016gop and Yb+ ion clock Banerjee:2023bjc ; and Optical cavities: H-maser comparison with a Si cavity (H/Si) Kennedy:2020bac ; while space-based current bound comes from Equivalence-principle tests: MICROSCOPE Berge:2017ovy . Future experiment projections are for the 229Th Nuclear Clock Antypas:2022asj and 171Yb+ ion clock Banerjee:2023bjc . Note: The Nuclear Clock line in previous estimates is cut off at $10^{-21}~{}\rm eV$, due to the Lyman-$\alpha$ constraint. However, there is no reason that the experiment cannot be run to search for lower mass range, as in $\rm Yb^{+}$ experiment. Therefore, we extrapolate linearly to lower masses. We are not aware of any space-based planned experiments that will constrain $d^{(2)}_{g}$. In cyan, we show the region of the parameter space critically screened by the Earth, as discussed in Sec. 2.2. Lyman-$\alpha$ Irsic:2017yje ; Rogers:2020ltq and ultra-faint dwarf galaxies (UFDs) Dalal:2022rmp also constrain ULDM. Finally, the regions of the parameter space corresponding to the black hole superradiance (BHSR) bounds Baryakhtar:2020gao ; Unal:2020jiy are displayed at the bottom of the plots by the arrows. In the terrestrial plots, we also show the region of parameter space critically screened by the Earth in cyan. See Sec. 2.2 for a detailed discussion of screening of quadratically-coupled scalars.
6 Conclusions